OpenStudy (anonymous):

5. If a baseball is projected upward from ground level with an initial velocity of 32 feet per second. a. What is the maximum height reached by the ball? b. How long is the ball airborne?

OpenStudy (anonymous):

what is the question? full question?

OpenStudy (anonymous):

5. If a baseball is projected upward from ground level with an initial velocity of 32 feet per second. a. What is the maximum height reached by the ball? b. How long is the ball airborne?

OpenStudy (anonymous):

Can somebody please help me? :)

OpenStudy (anonymous):

you will have to use equations of motion for this, have you read them?

OpenStudy (anonymous):

Is the equation -.5gt^2+Vot+So?

OpenStudy (ddcamp):

Yes, that's the right equation. What's V₀? S₀? Acceleration?

OpenStudy (anonymous):

there are mainly three of them.

OpenStudy (anonymous):

final velocity will be zero because you are throwing upwards.

OpenStudy (anonymous):

I believe Vo is initial velocity and So is initial starting point, but I am confused how to use the equation to solve

OpenStudy (anonymous):

acceleration will be equal to -g

OpenStudy (anonymous):

you know the value of g, right?

OpenStudy (anonymous):

I do not know the value of g, but I know that v is 32

OpenStudy (anonymous):

acceleration due to gravity, g = 9.8m/s^2

OpenStudy (ddcamp):

V_0 would be the initial velocity (the velocity the ball was thrown up with. S_0 would be 0 (the problem says we're on level ground) Acceleration due to gravity is -32 ft/s²

OpenStudy (anonymous):

it says we started from ground!!!

OpenStudy (anonymous):

then throw the ball from ground upwards because we are throwing upwards g which is acceleration due to gravity will be taken as negative

OpenStudy (anonymous):

Will g always be 9.8 in all velocity type problems? or will it sometimes be -16^2

OpenStudy (anonymous):

no, its a constant for only the problems involving acceleration due to gravity

OpenStudy (anonymous):

ok

OpenStudy (ddcamp):

g is 9.8 m/s² AND 32 ft/s². Different units. If the problem is in feet, use 32. If it's in meters, use 9.8/

OpenStudy (anonymous):

because g is acceleration, its unit is m/s^2 and unit of velocity is m/s

OpenStudy (ddcamp):

Ichiro, read the problem. We're not using metric! We're given feet, so we use feet. So now we have an equation: S(t) = 0 + 32t - 16t² Can you find the two times where S(t) = 0?

OpenStudy (anonymous):

2aS = final velocity ^2 - initial velocity ^2

OpenStudy (anonymous):

we dont use only 1 equation. and SI unit of length is meter so will have to convert first

OpenStudy (anonymous):

So the equation is -.5(9.8)t^2+32t+0 right?

OpenStudy (ddcamp):

We don't HAVE to use SI. It's good practice for actual science work, but we don't need to over complicate things... @monkeygirl8, you need to either use 32 ft/s² as g, or convert the initial velocity to m/s

OpenStudy (anonymous):

and we dont have to convert the feet in to meter here. initial velocity is 32feet per sec and final velocity is 0

OpenStudy (anonymous):

So the equation is -.5(9.8)t^2+32t+0 right?

OpenStudy (ddcamp):

No.

OpenStudy (anonymous):

nope convert the g or velocity

OpenStudy (anonymous):

write velocity in m/s or acceleration in feet/s^2

OpenStudy (anonymous):

Could someone please show me what the equation would look like, please :)

OpenStudy (anonymous):

wait a sec

OpenStudy (ddcamp):

S(t) = 0 + 32t - 16t² (this is in feet) OR S(t) = 0 + 9.7536t - 4.9t² (this is in meters) Remember, the problem is given in feet, so we need an answer in feet.

OpenStudy (ddcamp):

I would reccomend: S(t) = 0 + 32t - 16t²

OpenStudy (anonymous):

so I do not need to use g=9.8 anywhere?

OpenStudy (ddcamp):

No. That's only when the problem is given to you in meters.

OpenStudy (anonymous):

what would be the answer then, ddcamp?

OpenStudy (ddcamp):

The position is given by the equation: S(t) = c + bt + at² S(t) = 0 + 32t - 16t² We can find the maximum height by solving the equation: \[t=\frac{ -b }{ 2a } = \frac{ -32 }{ 2(-16) } = 1\] This gives us the time for the maximum point of the equation. Plug in 1 for t to find the height of the maximum.

OpenStudy (anonymous):

and answer?

OpenStudy (anonymous):

with my approx, it should be around 15 or 16 feet

OpenStudy (anonymous):

I am lost for what to do with the equation. I have a graphing calculator, can I just graph to find the maximum?

OpenStudy (anonymous):

\[2gS= Vf^2-Vi^2\] where g = 32.15 ft/s^2 Vf = 0 Vi = 32ft/s

OpenStudy (anonymous):

\[S=Vi^2/(2g)\]

OpenStudy (anonymous):

answer is 15.92ft

OpenStudy (anonymous):

So 15.92 is the max height. That makes sense. what is the amount of time it is airborne?

OpenStudy (anonymous):

now for the second part: \[Vf = V i+ g t\]

OpenStudy (ddcamp):

@monkeygirl8 What class is this for? Because you would use different methods for a math class and a physics class.

OpenStudy (anonymous):

\[Vf = 0\] \[g = 32.15 ft/s^2\] \[Vi = 32ft/s\]

OpenStudy (anonymous):

I am in pre calculus

OpenStudy (ddcamp):

OK. So, you're working with quadratics?

OpenStudy (anonymous):

yes

OpenStudy (ddcamp):

OK. So, we use the quadratic equation S(t) = 0 + 32t - 16t² (in math classes, we use g=32 ft/s² because that makes calculations easier) Can you find the zeroes of: S(t) = 0 + 32t - 16t² Then S(t)=0, we know the ball is at ground level.

OpenStudy (anonymous):

the zeros are 2 and 0

OpenStudy (ddcamp):

Good. So that means, at 2 seconds, the ball hits the ground again.

OpenStudy (anonymous):

Thank you so much! You were so much help :)

OpenStudy (ddcamp):

Did ichiros process for part (a) make sense? He used a physics formula, but I can explain in a way that uses pre-calc.

OpenStudy (anonymous):

I think I understand it now. Thank you so much for your help. :)

OpenStudy (anonymous):

hey ddcamp, which grade are you in?

OpenStudy (ddcamp):

I'm a senior. Taking linear algebra and college physics.